# Constraining Schwarzschild Models with Orbit Classifications

### v1

@@ Abstract
A simple orbit classification constraint extension to stellar dynamical modeling using Schwarzschild’s method is demonstrated. The classification scheme used is the existing ‘orbit circularity’ scheme ($\lambda_{z}$) where orbits are split into four groups - hot, warm, cold and counter rotating orbits. Other schemes which can be related to the orbit weights are expected to be viable as well. The results show that the classification constraint works well in modeling. However, given that orbits in external galaxies are not observable, it is not clear how the orbit classification for any particular galaxy may be determined. Perhaps range constraints for different types of galaxies determined from cosmological simulations may offer a way forward.
@@
galaxies: kinematics and dynamics – methods: numerical
@@ 1 Introduction
The stellar dynamics of galaxies may be investigated by creating models of galaxies, constraining them with observational data, and then examining the models to see what might be learnt about the real galaxies. Binney & Tremaine (2008) describes the theory behind some of the traditional modeling techniques that might be used. A key point which must be accepted about external galaxy models is that the models can only be indicative or illustrative of the galaxies because of the current technical limitations of the instruments employed for observations: six dimensional (6D) galaxy phase space data are not obtainable. Mathematically, recovery of the phase space data from observations is not possible in general. Related deprojection issues are included, for example, in Rybicki (1987), and, more recently, in papers by Cappellari (2020) and Vasiliev & Valluri (2020). Long et al. (2021) demonstrated that Schwarzschild’s method does not comply fully with the theoretical statements in Rybicki (1987). Observing and modeling our Galaxy is different in that the instrumental capability to obtain 3D spatial and velocity data is available. As a consequence, for example, it is not clear to what extent orbits in galaxies in cosmological simulations are truly representative of those present in real galaxies. Specialized n-body simulations help us gain insight into the orbits that might comprise particular components of galaxies, e.g. galactic bars, but do not give us the capability to model all aspects of real galaxies accurately.
It has been recognized that applying machine learning (ML) modeling methods to physical systems requires that the relevant physical laws should be adhered, and has resulted in, for example, the development of physics informed neural networks (Raissi et al., 2019). It is appropriate to revisit more traditional stellar dynamical modeling methods, such as Schwarzschild’s method (Schwarzschild, 1979) or Syer & Tremaine’s made-to-measure method (Syer & Tremaine, 1996), to understand whether their solutions are astrophysically realistic or not. In this investigation, Schwarzschild’s method which superimposes weighted stellar orbits to try and reproduce observational data is examined.
@@
    Low orbit utilization (without regularization) in that only $\approx 10\%$ of orbits supplied are an effective part of any solution - see table 2 of Jin et al. (2019). This percentage becomes even lower if chemo-dynamical techniques such as in Long & Mao (2018) are applied. Regularization can be used to improve the utilization to typically $\approx 75\%$ or more depending on the optimization method.
@@
Given that Schwarzschild’s method can not be used without amendment or enhancement to determine the overall orbit classification for a galaxy, if the orbit classification were to be known from other sources, could that classification be included in Schwarzschild modeling as a constraint perhaps ? Answering this question is the focus of this article. The article’s structure is as follows. Section 2 describes at a top level the approach taken in the investigations. The galaxy data used is described in Section 3, and the relevant theory and descriptions of the methods are in Section 4. Results and subsequent discussion are in Sections 5 and 6, with conclusions in Section 7.
@@ 2 Approach
Whereas the Introduction, Section 1, sets the context and describes what will be investigated, this section describes how how it will be done. A number of axisymmetric models of galaxies are constructed, under varying conditions, using an existing software implementation of Schwarzschild’s method taken from Long et al. (2021), and analyzed. Line-of-sight luminosity and kinematical data from nine galaxies in total are used, with four real galaxies from ATLAS3D (Cappellari et al., 2011) and five simulated galaxies from Illustris TNG (Nelson et al., 2019). The simulated galaxies have known orbit circularity classifications (Xu et al., 2019).
@@
Modeling runs are accomplished in 3 stages:
@@ 3 Galaxy Data
For the ATLAS3D galaxies (NGC 1248, NGC 3838, NGC 4452, and NGC 4551), the data have been produced by the ATLAS3D survey (Cappellari et al., 2011). The four galaxies are the same galaxies as used in Long (2016) and Long & Mao (2018), and have various inclinations, different mass-to-light ratios, and sense of bulk rotation (see Tab. 1). Surface brightness (SB) values are calculated, using multi-Gaussian expansions (Sect. 4.1), on a $(16,16)$ polar grid with equal interval radial distances, and no straddling of a galaxy’s major axis. The IFU cells for kinematical data (first and second velocity moments) are the Voronoi cells resulting from signal-to-noise processing of the raw data.
@@
The inclinations are taken from Cappellari et al. (2013), while the mass-to-light ratios were determined by M2M modeling in Long (2016). The SB maximum radius gives the overall size of the polar grid used for SB data values. The IFU (integral field unit) column gives the number of Voronoi cells for the kinematical data. The numbers in brackets are the maximum ‘on sky’ X and Y values from the galaxy’s cell centroids, and give a rough indication of the kinematical data coverage without needing to refer to the convex hull surrounding the cell centroids.
The same Illustris TNG galaxies (Nelson et al., 2019) as in Long et al. (2021) are used. The selection strategy for the galaxies is described in Long et al. (2021), and is not repeated here. The five axisymmetric galaxies (with identifiers, A490, A1090, A1190, A1290, and A1390) are viewed edge on with the mass-to-light ratio taken as $1.0$ for all galaxies. It has been assumed that stellar particles kinematics can be validly be used to create stellar kinematical data as if from a real galaxy (consistent with Li et al. (2016) or Xu et al. (2019), for example). Surface brightness (SB) values are calculated, using multi-Gaussian expansions, on a $(16,16)$ polar grid with logarithmic interval radial distances, and straddling the galaxy’s major axis. The IFU cells for kinematical data (the Gauss-Hermite coefficients $h_{1}$ to $h_{4}$) are constructed to mimic a real galaxy’s data. Table 2 is the equivalent of Table 1 for the simulated galaxies.
@@
As in Table 1, the SB maximum radius gives the overall size of the SB polar grid, and the IFU column gives the number of Voronoi cells with the numbers in brackets being the maximum X and Y values from the galaxy’s cell centroids. Inclinations for all the simulated galaxies are $90^{\circ}$, and mass-to-light ratios are $1.0$.
The IFU cells for the ATLAS3D galaxies and the Illustris simulated galaxies contain the observational kinematical data that Schwarzschild’s method is asked to replicate. Data are symmetrized appropriately for an axisymmetric model/potential. The design matrix size used in the convex optimization (see equation 1) may be calculated as the total number of data points for observables times the number of orbits. As an example, NGC 3838 has surface brightness (256 points) and two IFU observables (2 x 383 = 766 points) giving 1022 data points in total. For 8000 orbits, the design matrix size is 1022 rows and 8000 columns. For simulated galaxy A1190 with four IFU observables, the matrix size 696 rows and 8000 columns.
@@ 4.1 Gravitational Potentials
Galaxy gravitational potentials and derivatives are derived from multi-Gaussian expansions (MGEs) of the galaxies’ surface brightness (Emsellem et al., 1994; Cappellari, 2002). For computer performance reasons, an MGE potential and its gradients are implemented as a set of interpolation tables, for example as in van den Bosch et al. (2008) or Long & Mao (2012).
@@ 4.2 Orbit Initial Conditions
The initial phase space coordinates for orbits are created using the three integral space based scheme (3I). Three variants are created by varying the sense of rotation of the orbits:
@@
    as in (2) but with 100% matching the data’s sense of rotation and no counter rotating orbits.
Similarly to Zhu et al. (2018), the circularity measure $\lambda_{z}$, the ratio of the orbit’s $L_{z}$ component of angular momentum to the maximum value allowed by the orbit’s energy, is used to classify orbits into one of hot ($\left|\lambda_{z}\right|<=0.25$), cold ($\lambda_{z}>=0.80)$, warm ($0.80.25<\lambda_{z}<$ 0.8) or counter-rotating (the remainder). Practically, internally, it is more convenient to split the orbits into six classes - rotating or counter-rotating and then one of hot, warm or cold - and only combine classes where necessary for presentational or comparison purposes. These classes are used extensively in analyzing modeling runs. The total orbit weights in each of the classes are taken as the overall orbit classification for a galaxy.
Note that the combination of a steady state gravitational potential for a galaxy and the initial spatial and velocity conditions for an orbit mean that the orbit’s trajectory is fully determined. From Hamiltonian mechanics, energy is a conserved quantity for an orbit, and, in addition, for an axisymmetric galaxy potential so is the component of angular momentum ($L_{z}$) parallel to the symmetry axis of the galaxy. This means that the $\lambda_{z}$ classifier of orbits as hot, warm, or cold can be found directly from the initial conditions. As a consequence, if all the orbits in a model are regarded as active orbits (not zero-weighted) by the optimization method then the relative proportions of the numbers of hot, warm and cold orbits must match the proportions from the initial conditions. This remains true even for gravitational systems where $L_{z}$ is not conserved but some time or orbit averaged value is used instead. What is of more interest is the relative proportions of the orbit weights allocated to hot, warm and cold orbits as these proportions can be related to the global properties of a galaxy.
@@ 4.4 Constraints using Orbit Classifications
In this section, the weighting process is first considered in optimization objective function terms (rather matrices). Usually the Schwarzschild objective function is framed in terms of reproducing the observed luminosity and kinematical data so the optimizing process selects and weights orbits to try and cause this reproduction to happen. If the objective function is framed to achieve a different end effect (for example, related to the distribution of some other quantity), a set of orbits appropriate to that objective would be selected and weighted. What is needed here is that not only are observables reproduced but that some astrophysically realistic mix of orbit types is used in so doing.
@@
Structurally, expression (3) has the same $\chi^{2}$ form as the observable data in the objective function. In this article, expression (2) is used, together with an orbit classification parameter $\mu_{\rm{oc}}$ to set the overall importance of the expression amongst all the other terms in the function. Here, $\mu_{\rm{oc}}=10^{3}$.
Reverting now to matrix terms, each class of orbit type $C_{j}$ is represented in matrix $\mathbf{D}$ (see equation 1) by a row with each column entry having value $1$ if orbit $i$ is in $C_{j}$ and 00 otherwise. The corresponding entries in matrix $\mathbf{K}$ are the $T_{j}$ values. In the context of $\lambda_{z}$ circularity classifications, the [red, warm, cold, counter-rotating] target values might, for example, be [30%, 45%, 15%, 10%] with 4 matrix rows being required, and the entries per row indicating whether the individual orbits are of the associated type (entry value = 1) or not (entry value = 0). Note that the mechanism described is not specific to $\lambda_{z}$ classifications but could be used with other classification schemes as well. Also, the approach described here is readily adaptable to the made-to-measure method.
@@ 4.5 Regularization
Smoothing the weight distribution is performed in integral space and is described below. For axisymmetric models, because of uncertainties about quite what form the third integral takes, only the conserved quantities energy ($E$) and the $L_{z}$ component of angular momentum are used. Working with the initial conditions on the orbits, each orbit’s nearest two neighbors in $(E,L_{z})$ space are determined, and the second derivatives of the weights using finite difference formulae are calculated at the orbit’s $(E,L_{z})$ position. A two dimensional tree is used to find the nearest neighbors and the finite difference formula used (backward, forward or central) depends on the relative positioning of the orbit and its neighbors. Solution or curvature smoothing is a well-established technique, see Boyd & Vandenberghe (2004) or Press et al. (2007), for example.
The regularization parameters are $1.6\times 10^{3}$ for penalization, $6.4\times 10^{2}$ for smoothing, and were determined by experimentation. For convenience, the same values are used for all galaxies. Note that both forms of regularization can be used individually in isolation or combined together to give a composite regularization term. For the remainder of this article, the terms regularization (for heavy weights penalization) and smoothing (for second derivative smoothing) are employed.
@@ 5.1 Baseline Models
A set of models are constructed and analyzed to establish a baseline of understanding, or a reference point, for subsequent modeling. Both sets of galaxies are used for this task. 3I initial conditions, with the rotation of 75% of the orbits aligned to the bulk rotation of the galaxies, are employed. The orbit classification resulting from the weighted orbits is calculated and, for the simulated galaxies, compared with the classification in Xu et al. (2019).
@@
Looking at the mean $\chi^{2}$ column, the observed data values are reproduced well by the modeling. For models without regularization, mean $\chi^{2}<<1$ indicate that significant overfitting is taking place. The heaviest weighted orbits for a galaxy are those orbits remaining close (in projection) to the galaxy center with the lighter weighted orbits having larger apocenters. Figure 1 shows this weight progression for galaxy A1190 as an example. The position of the heaviest weights centrally matches the region where the surface brightness observable dominates numerically other observables in the model.
@@
Using a simple circularity classification into hot, warm and cold orbits with no separation by sense of rotation, it is clear that orbits in different classes tend to have a role to play in particular regions - see Figure 2, this time using galaxy A1390 as an example.
@@
As regularization is penalizing heavy weights, the maximum orbit weights are reduced (by approximately an order of magnitude), and the weight ranges are extended to lower values. Regularization does cause more orbits to be active. However the issue concerning a small number of orbits being allocated heavy weights remains: it has just been ‘down-shifted’ by a factor of 10 - see the top row of Figure 3.
@@
Smoothing alone also causes more orbits to be active but some orbits have even heavier weights as the orbit count distribution is broadened overall, generally to lower weights - see the bottom row of Figure 3. Combining regularization and smoothing together produces a result more akin to regularization alone but with more active orbits, and a performance penalty (see Tab. 8).

### v2

@@ Abstract
A simple orbit classification constraint extension to stellar dynamical modeling using Schwarzschild’s method is demonstrated. The classification scheme used is the existing ‘orbit circularity’ scheme ($\lambda_{z}$) where orbits are split into four groups—hot, warm, cold and counter-rotating orbits. Other schemes which can be related to the orbit weights are expected to be viable as well. The results show that the classification constraint works well in modeling. However, given that orbits in external galaxies are not observable, it is not clear how the orbit classification for any particular galaxy may be determined. Perhaps range constraints for different types of galaxies determined from cosmological simulations may offer a way forward.
@@
galaxies: kinematics and dynamics, galaxies: general, methods: numerical
@@ 1 Introduction
The stellar dynamics of galaxies may be investigated by creating models of galaxies, constraining them with observational data, and then examining the models to see what might be learnt about the real galaxies. The book Binney & Tremaine (2008) describes the theory behind some of the traditional modeling techniques that might be used. A key point which must be accepted about external galaxy models is that the models can only be indicative or illustrative of the galaxies because of the current technical limitations of the instruments employed for observations: six-dimensional (6D) galaxy phase space data are not obtainable. Mathematically, recovery of the phase space data from observations is not possible in general. Related deprojection issues are included, for example, in Rybicki (1987), and, more recently, in papers by Cappellari (2020) and Vasiliev & Valluri (2020). Long et al. (2021) demonstrated that Schwarzschild’s method does not comply fully with the theoretical statements in Rybicki (1987). Observing and modeling our Galaxy are different in that the instrumental capability to obtain three-dimensional (3D) spatial and velocity data is available. As a consequence, for example, it is not clear to what extent orbits in galaxies in cosmological simulations are truly representative of those present in real galaxies. Specialized n-body simulations help us gain insight into the orbits that might comprise particular components of galaxies, e.g. galactic bars, but do not give us the capability to model all aspects of real galaxies accurately.
It has been recognized that applying machine learning modeling methods to physical systems requires that the relevant physical laws should be adhered, and has resulted in, for example, the development of physics informed neural networks (Raissi et al., 2019). It is appropriate to revisit more traditional stellar dynamical modeling methods, such as Schwarzschild’s method (Schwarzschild, 1979) or Syer & Tremaine’s made-to-measure method (Syer & Tremaine, 1996), to understand whether their solutions are astrophysically realistic or not. In this investigation, Schwarzschild’s method which superimposes weighted stellar orbits to try and reproduce observational data is examined.
@@
    Low orbit utilization (without regularization) in that only $\approx 10\%$ of orbits supplied are an effective part of any solution—see Table 2 of Jin et al. (2019). This percentage becomes even lower if chemo-dynamical techniques such as in Long & Mao (2018) are applied. Regularization can be used to improve the utilization to typically $\approx 75\%$ or more depending on the optimization method.
@@
Given that Schwarzschild’s method cannot be used without amendment or enhancement to determine the overall orbit classification for a galaxy, if the orbit classification were to be known from other sources, could that classification be included in Schwarzschild modeling as a constraint perhaps? Answering this question is the focus of this article. The article’s structure is as follows. Section 2 describes at a top level the approach taken in the investigations. The galaxy data used are described in Section 3, and the relevant theory and descriptions of the methods are in Section 4. Results and subsequent discussion are in Sections 5 and 6, with conclusions in Section 7.
@@ 2 Approach
Whereas the Introduction, Section 1, sets the context and describes what will be investigated, this section describes how it will be done. A number of axisymmetric models of galaxies are constructed, under varying conditions, using an existing software implementation of Schwarzschild’s method taken from Long et al. (2021), and analyzed. Line-of-sight luminosity and kinematical data from nine galaxies in total are used, with four real galaxies from ATLAS3D (Cappellari et al., 2011) and five simulated galaxies from IllustrisTNG (Nelson et al., 2019). The simulated galaxies have known orbit circularity classifications (Xu et al., 2019).
@@
Modeling runs are accomplished in three stages:
@@ 3 Galaxy Data
For the ATLAS3D galaxies (NGC 1248, NGC 3838, NGC 4452, and NGC 4551), the data have been produced by the ATLAS3D survey (Cappellari et al., 2011). The four galaxies are the same galaxies as used in Long (2016) and Long & Mao (2018), and have various inclinations, different mass-to-light ratios, and sense of bulk rotation (see Table 1). Surface brightness values are calculated, using multi-Gaussian expansions (MGEs) (Sect. 4.1), on a $(16,16)$ polar grid with equal interval radial distances, and no straddling of a galaxy’s major axis. The integral field unit (IFU) cells for kinematical data (first and second velocity moments) are the Voronoi cells resulting from signal-to-noise processing of the raw data.
@@
The inclinations are taken from Cappellari et al. (2013), while the mass-to-light ratios (in the M/L Ratio column) were determined by made-to-measure modeling in Long (2016). The SB maximum radius column gives the overall size of the polar grid used for surface brightness data values. The IFU column gives the number of Voronoi cells for the kinematical data. The numbers in brackets are the maximum ‘on sky’ X and Y values from the galaxy’s cell centroids, and give a rough indication of the kinematical data coverage without needing to refer to the convex hull surrounding the cell centroids.
The same IllustrisTNG galaxies (Nelson et al., 2019) as in Long et al. (2021) are used. The selection strategy for the galaxies is described in Long et al. (2021), and is not repeated here. The five axisymmetric galaxies (with identifiers, A490, A1090, A1190, A1290, and A1390) are viewed edge on with the mass-to-light ratio taken as $1.0$ for all galaxies. It has been assumed that the kinematics of stellar particles can validly be used to create stellar kinematical data as if from a real galaxy (consistent with Li et al. (2016) or Xu et al. (2019), for example). Surface brightness values are calculated, using MGEs, on a $(16,16)$ polar grid with logarithmic interval radial distances, and straddling the galaxy’s major axis. The IFU cells for kinematical data (the Gauss-Hermite coefficients $h_{1}$ to $h_{4}$) are constructed to mimic a real galaxy’s data. Table 2 is the equivalent of Table 1 for the simulated galaxies.
@@
As in Table 1, the SB maximum radius column gives the overall size of the surface brightness polar grid, and the IFU column gives the number of Voronoi cells with the numbers in brackets being the maximum X and Y values from the galaxy’s cell centroids. Inclinations for all the simulated galaxies are $90^{\circ}$, and mass-to-light ratios are $1.0$.
The IFU cells for the ATLAS3D galaxies and the IllustrisTNG simulated galaxies contain the observational kinematical data that Schwarzschild’s method is asked to replicate. Data are symmetrized appropriately for an axisymmetric model/potential. The design matrix size used in the convex optimization (see equation 1) may be calculated as the total number of data points for observables times the number of orbits. As an example, NGC 3838 has surface brightness (256 points) and two IFU observables (2 x 383 = 766 points) giving 1022 data points in total. For 8000 orbits, the design matrix size is 1022 rows and 8000 columns. For simulated galaxy A1190 with four IFU observables, the matrix size is 696 rows and 8000 columns.
@@ 4.1 Gravitational Potentials
Galaxy gravitational potentials and derivatives are derived from MGEs of the galaxies’ surface brightness (Emsellem et al., 1994; Cappellari, 2002). For computer performance reasons, an MGE potential and its gradients are implemented as a set of interpolation tables, for example as in van den Bosch et al. (2008) or Long & Mao (2012).
@@ 4.2 Orbit Initial Conditions
The initial phase space coordinates for orbits are created using the 3I scheme. Three variants are created by varying the sense of rotation of the orbits:
@@
    as in (2) but with 100% matching the data’s sense of rotation and no counter-rotating orbits.
Similarly to Zhu et al. (2018), the circularity measure $\lambda_{z}$, the ratio of the orbit’s $L_{z}$ component of angular momentum to the maximum value allowed by the orbit’s energy, is used to classify orbits into one of hot ($\left|\lambda_{z}\right|<=0.25$), cold ($\lambda_{z}>=0.80)$, warm ($0.80.25<\lambda_{z}<$ 0.8) or counter-rotating (the remainder). Practically, internally, it is more convenient to split the orbits into six classes—rotating or counter-rotating and then one of hot, warm or cold—and only combine classes where necessary for presentational or comparison purposes. These classes are used extensively in analyzing modeling runs. The total orbit weights in each of the classes are taken as the overall orbit classification for a galaxy.
Note that the combination of a steady state gravitational potential for a galaxy and the initial spatial and velocity conditions for an orbit means that the orbit’s trajectory is fully determined. From Hamiltonian mechanics, energy is a conserved quantity for an orbit, and, in addition, for an axisymmetric galaxy potential so is the component of angular momentum ($L_{z}$) parallel to the symmetry axis of the galaxy. This means that the $\lambda_{z}$ classifier of orbits as hot, warm, or cold can be found directly from the initial conditions. As a consequence, if all the orbits in a model are regarded as active orbits (not zero-weighted) by the optimization method then the relative proportions of the numbers of hot, warm and cold orbits must match the proportions from the initial conditions. This remains true even for gravitational systems where $L_{z}$ is not conserved but some time or orbit averaged value is used instead. What is of more interest is the relative proportions of the orbit weights allocated to hot, warm and cold orbits as these proportions can be related to the global properties of a galaxy.
@@ 4.4 Constraints using Orbit Classifications
In this section, the weighting process is first considered in optimization objective function terms (rather than matrices). Usually the Schwarzschild objective function is framed in terms of reproducing the observed luminosity and kinematical data so the optimizing process selects and weights orbits to try and cause this reproduction to happen. If the objective function is framed to achieve a different end effect (for example, related to the distribution of some other quantity), a set of orbits appropriate to that objective would be selected and weighted. What is needed here is that not only are observables reproduced but that some astrophysically realistic mix of orbit types is used in so doing.
@@
Structurally, expression (3) has the same $\chi^{2}$ form as the observable data in the objective function. In this article, expression (2) is used, together with an orbit classification parameter $\mu_{\rm{oc}}$, to set the overall importance of the expression amongst all the other terms in the function. Here, $\mu_{\rm{oc}}=10^{3}$.
Reverting now to matrix terms, each class of orbit type $C_{j}$ is represented in matrix $\mathbf{D}$ (see equation 1) by a row with each column entry having value $1$ if orbit $i$ is in $C_{j}$ and 00 otherwise. The corresponding entries in matrix $\mathbf{K}$ are the $T_{j}$ values. In the context of $\lambda_{z}$ circularity classifications, the [red, warm, cold, counter-rotating] target values might, for example, be [30%, 45%, 15%, 10%] with four matrix rows being required, and the entries per row indicating whether the individual orbits are of the associated type (entry value = 1) or not (entry value = 0). Note that the mechanism described is not specific to $\lambda_{z}$ classifications but could be used with other classification schemes as well. Also, the approach described here is readily adaptable to the made-to-measure method.
@@ 4.5 Regularization
Smoothing the weight distribution is performed in integral space and is described below. For axisymmetric models, because of uncertainties about quite what form the third integral takes, only the conserved quantities energy ($E$) and the $L_{z}$ component of angular momentum are used. Working with the initial conditions on the orbits, each orbit’s nearest two neighbors in $(E,L_{z})$ space are determined, and the second derivatives of the weights using finite difference formulae are calculated at the orbit’s $(E,L_{z})$ position. A two-dimensional tree is used to find the nearest neighbors and the finite difference formula used (backward, forward or central) depends on the relative positioning of the orbit and its neighbors. Solution or curvature smoothing is a well-established technique, see Boyd & Vandenberghe (2004) or Press et al. (2007), for example.
The regularization parameters are $1.6\times 10^{3}$ for penalization and $6.4\times 10^{2}$ for smoothing, and were determined by experimentation. For convenience, the same values are used for all galaxies. Note that both forms of regularization can be used individually in isolation or combined together to give a composite regularization term. For the remainder of this article, the terms regularization (for heavy weights penalization) and smoothing (for second derivative smoothing) are employed.
@@ 5.1 Baseline Models
A set of models is constructed and analyzed to establish a baseline of understanding, or a reference point, for subsequent modeling. Both sets of galaxies are used for this task. 3I initial conditions, with the rotation of 75% of the orbits aligned to the bulk rotation of the galaxies, are employed. The orbit classification resulting from the weighted orbits is calculated and, for the simulated galaxies, compared with the classification in Xu et al. (2019).
@@
Looking at the mean $\chi^{2}$ column, the observed data values are reproduced well by the modeling. For models without regularization, mean $\chi^{2}$ values $<<1$ indicate that significant overfitting is taking place. The heaviest weighted orbits for a galaxy are those orbits remaining close (in projection) to the galaxy center with the lighter weighted orbits having larger apocenters. Figure 1 shows this weight progression for galaxy A1190 as an example. The position of the heaviest weights centrally matches the region where the surface brightness observable dominates numerically other observables in the model.
@@
Using a simple circularity classification into hot, warm and cold orbits with no separation by sense of rotation, it is clear that orbits in different classes tend to have a role to play in particular regions—see Figure 2, this time using galaxy A1390 as an example.
@@
As regularization is penalizing heavy weights, the maximum orbit weights are reduced (by approximately an order of magnitude), and the weight ranges are extended to lower values. Regularization does cause more orbits to be active. However the issue concerning a small number of orbits being allocated heavy weights remains: it has just been ‘down-shifted’ by a factor of 10—see the top row of Figure 3.
@@
Smoothing alone also causes more orbits to be active but some orbits have even heavier weights as the orbit count distribution is broadened overall, generally to lower weights—see the bottom row of Figure 3. Combining regularization and smoothing together produces a result more akin to regularization alone but with more active orbits, and a performance penalty (see Table 8).